How do u solve the equation: 3p-2(p-4)= 7p+6 ....wat does P equal.....iem very confused with this problem and i dont really understand how 2 do it!
Shelly~

Answer by jessicawhidden(3)

follow the order of operations, don't allow yourself to be confused by the variable, it is just an unknown number that you are trying to find.
3p-2(p-4)=7p+6---multiply to get ride of the ( )pay attention to the - sign.
-2p--8
3p-2p+8=7p+6-----combine like terms.
3p-2p
p+8=7p+6---------get like parties to one side of the equation.
-p from both sides of the = sign
8=6p+6
-6 from both sides of the = sign
2=6p
divide both sides by 6
2/6=p
reduce your fraction
p=1/3
you can plug your answer into the problem to check.
3(1/3)-2(1/3-4)=7(1/3)+6
3/3-2/3+8=7/3+6
multiply both sides by 3 to get ride of your fractions
3-2+24=7+18
solve
25=25